OpenStudy (anonymous):

The English Alphabet has 20 letters more than one−fourth the number of letters in the Greek Alphabet. The English Alphabet has 26 letters. Which of the following equations would you use to find the number of letters, g, in the Greek Alphabet

OpenStudy (anonymous):

"which of the following equations" do yu have a list to choose from?

OpenStudy (anonymous):

The sum of g and 20 all over 4 equals 26 4g − 20 = 26 1 over 4g + 20 = 26 4g − 26 = 20

OpenStudy (anonymous):

i think its the 3rd one

OpenStudy (anonymous):

ok